Circular No. 8240

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2002 LZ_{11} (LINEAR) An apparently asteroidal object with not-unusual motion recorded by

the LINEAR project (discovery observation from MPS 57856 below), and linked over 2002 June 5–13 and designated by the Minor Planet Center as 2002 LZ_{11} (but indicated on MPEC 2002-L64 as having a cometary orbit), was accidentally reobserved by LINEAR on numerous occasions in July 2002 and during July–Nov. 2003 (and by LONEOS on 2003 Oct. 16). E. Christensen, University of Arizona, recognized 2002 LZ_{11} as having cometary appearance yesterday in CCD images (mag 16.6–16.9) obtained accidentally by the Catalina Sky Survey (0.68-m f/1.8 Schmidt) on 2003 Oct. 29. Further images yesterday by Christensen apparently confirmed the cometary appearance, as did images obtained today at the request of the Minor Planet Center by J. Young (0.61-m reflector at Table Mountain Observatory), who described the object as having a round coma, 7′′ across, with a featureless fanshaped tail ∼ 15′′–20′′ long in p.a. 190^{o}–280^{o}.

2002 UT  α_{2000}  δ_{2000}  Mag.
June 5.29966  17^{h}47^{m}23.^{s}52  −14^{o}14′39′′.8  18.1

The following elliptical orbital elements by B. G. Marsden are from 69 observations (all published in MPSs) made during 2002 June 5–2003 Nov. 2 (mean residual 0′′.7). The ephemeris magnitudes assume H_{10} = 11.0.

Epoch = 2003 Mar. 22.0 TT

T = 2003 Mar. 15.6630 TT  ω = 107.^{o}9740
e = 0.351722  Ω = 231.1424 }2000.0
q = 2.370483 AU  i = 11.5194

a = 3.656582 AU n^{o} = 0.1409584 P = 6.99 years

2003/04 α_{2000}  δ_{2000}  Δ  r  ϵ  β  Mag.
Nov. 7 2^{h}50.^{m}85 +18^{o}21.^{′}1 1.812 2.802 177.^{o}5 0.^{o}9 16.8
17 2 43.58 +17 05.0 1.855 2.833 169.5  3.6 16.9
27 2 37.47 +15 54.3 1.927 2.865 157.7  7.5 17.0
Dec. 7 2 33.19 +14 54.6 2.025 2.897 146.3 10.9 17.2
17 2 31.06 +14 09.0 2.146 2.929 135.4 13.6 17.3
27 2 31.19 +13 38.6 2.286 2.962 125.1 15.8 17.5
Jan. 6 2 33.50 +13 23.0 2.440 2.995 115.3 17.3 17.7
16 2 37.80 +13 20.5 2.604 3.028 106.1 18.2 17.9
26 2 43.86 +13 29.0 2.775 3.062  97.4 18.6 18.1
2003 November 18  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green